# Visual Analysis of Multi-outcome Causal Graphs

### v1

@@ Abstract
The overall process of the visual analysis of multi-outcome causal graphs of a health research data [15]. Analysts use (1) the Single Graph View to analyze and edit the causal graph (a–c) of a single outcome aided by a set of new visualization techniques (d–f) to leverage the synergy of various causal discovery methods. The fine-tuned causal graphs of single outcomes are stored in (2) the History View (g). Next, causal graphs of interest are selected for visual comparison in (3) the Multi-outcome Graphs Comparison View (h, i) supported by our new graph layout and visual mappings (i).
@@ Introduction
In medicine, experts also seek to discover causalities between potential risk factors and a specific disease, aiming to explore factors influencing the development of a particular disease and eventually intervene in the prognosis. Except for the typical case of studying the causality of a single disease (a single outcome variable), understanding the similarities and differences between different yet related causal relationships among different diseases, i.e., multiple outcome variables, is also important. For example, a study into the co-occurrence of more than one chronic condition in one person—multimorbidity [28], and the combined effect of multiple conditions in one person, i.e., comorbidity [60], is beneficial for proper interventions and treatments. Therefore, it is necessary to perform analysis and comparisons for causality of multiple outcomes. Unfortunately, most previous visual analysis works intend to address problems with only one particular outcome.
In this paper, we study the causalities of multiple outcome variables that have not been considered in visual analytics. Based on a requirement analysis in close collaboration with medical experts, we propose a visual analysis framework as shown in Visual Analysis of Multi-outcome Causal Graphs. The framework consists of two stages: the discovery and interactive fine-tuning of a causal graph of a single outcome (Visual Analysis of Multi-outcome Causal Graphs(1)); and the comparative analysis of causal graphs of multiple outcomes (Visual Analysis of Multi-outcome Causal Graphs(3)).
Causal graphs in the first stage are calculated using three causal discovery algorithms: a classic causal discovery method that provides a basic graph, and two state-of-the-art methods are applied to provide analysts with more suggestions. The resulting graphs are visualized progressively as node-link diagrams with a superposed comparative visual design for easy detection of differences. The underlying machine learning methods can be interactively controlled, e.g., pausing and resuming, to reduce computation time. Users then interactively fine-tune the single outcome causal graph with reference to multiple views of multivariate data visualizations and with their domain knowledge.
@@
We evaluate our method with numerical measurements, a case study, and an expert user study with medical experts. The effectiveness of the analysis and visualization of single outcome graphs is demonstrated using benchmark datasets; the multi-outcome comparison design is evaluated with the stress measurement. The overall visual analysis method is evaluated through a case study of the analysis of diseases that are usually seen in multimorbidity/comorbidity using two widely accepted health research datasets. Feedback from expert users indicates that our method is useful for supporting causal reasoning of single and multiple outcomes and provides insights with the comparison of multi-outcome causal graphs that are impossible with existing methods.
@@
Please note that our method aims to address a domain problem in healthcare but is also general enough to aid causal analysis of single and multiple outcomes. The source code of our visual analysis system is publicly available at: https://github.com/mengjiefan/multi_outcome/tree/vis_rev_sub.
@@ 1 Related Work
Our work is closely related to three areas of research: causal discovery, visualization and visual analysis of causality, and comparative visualization.
@@ 1.1 Causal Discovery
Combinatoric methods search for DAGs satisfying criteria within the aforementioned four categories. Popular combinatoric methods include the Peter-Clark (PC) algorithm [55], GES [16], satisfiability (SAT)-solver [32], and LinGAM [53].
In contrast, continuous methods rely on continuous optimizations that are typically used in neural networks. NO TEARS (Non-combinatoric Optimization via Trace Exponential Augmented Lagrangian Structure learning) is considered the first to formulate the combinatorial graph search method as a continuous optimization problem [70]. Further improvement of the method that is faster and more accurate than NO TEARS is available [9]. DAG-GNN [67] is a graph neural network approach to solve the continuous optimization problem. Methods focused on categorical data are also available with continuous optimizations using interventional or reinforcement learning [37, 12].
One issue with causal discovery methods is that they typically assume that all variables have the same data type–either categorical or continuous. In real-world datasets, however, data types are mixed, for example, typical survey data in medicine includes continuous biological measures like body height, weight, and blood pressure, and binary/categorical attributes like gender and stages of diseases. Although few but dedicated methods are available for mixed-type datasets [3, 4, 40, 14], while some others, for example, DAG-GNN [67], can handle such data inherently. A latent-PC algorithm is devised for multi-dimensional mixed-type data that is typical in medical research [14].
@@ 1.2 Visualization of Causal Graphs and Visual Analysis of Causal Reasoning
The visualization of causal graphs is studied in several works [6, 66, 61, 19]. Different visual mappings of the graph, e.g., the adjacent matrix, and node-link diagram are compared in a user study [61]. For node-link diagrams, different graph layouts are also compared with user studies [6, 61]. Layered graph layouts are used for showing large causal graphs [66] or comparing different causal discovery results [19].
Visual analysis methods for causal reasoning are available to aid the understanding of causalities and support interactive decision-making. An interactive causal reasoning interface is introduced to visually edit, and verify causal relationships with multiple linked views of the causal graph, and statistical parameters [63]. A more comprehensive extension of the method is available with the support of multidimensional data visualizations, a visual design of the causal graph with layered graph layout, and the capability of handling mixed-type data [64]. An exploratory causal analysis approach introduces the uncertainty-aware visualization of causal graphs and provides interactive means for exploring, validating, and applying causal relations for decision-making [66].
Some works focus on the understanding of causal inferences and counterfactuals [38]. A multi-view visualization tool is available to support iterative causal inference [26]. Counterfactuals are used to improve the visualization of relationships between variables [36] and causal inferences [10]. Visual analytics of time-dependent causality is devised for event sequences with a user-modifiable Granger causality model [34]. Causal hypotheses of time-dependent events can be formulated and tested with time delay in a visual analysis method [65].
@@ 1.3 Comparative Visualization
It is a common yet challenging task to design methods supporting comparison in visualization [24]. Four considerations are presented for comparative visualization designing and evaluation [24]. A general taxonomy is devised for visual comparison classifying visual designs into three basic categories—juxtaposition, superposition, and explicit encodings [25]. These categories should be combined for effective comparative visualization tools and tasks. A qualitative user study shows that interactive techniques are important to visual comparison tasks [58]. A combined encoding of colors and positions enables analysts to easily view differences and changes in the comparison of task-driven topic models [1].
Research efforts are made for the visual comparison of graphs. Visual comparison of two graphs is achieved by visualizing their similarities and differences in a merged graph [5]. Interactive graph matching [29] addresses the visual comparison of graphs and clustered graphs. Weighted graph comparison is explored in brain connectivity analysis [2] where the effectiveness of node-link diagrams and adjacency matrices are compared. A map-based abstraction of node-aligned graphs on a triangle mesh facilitates a clutter-reduced comparison of undirected graphs [33]. Layout optimization, graph embedding, and clustering and grouping algorithms are used for the comparison of two temporal graph data [69].
@@ 2 Requirement Analysis and Method Overview
Our work is a long-term collaboration with medical experts, which started in 2022. The work was initially motivated by the needs of one of our collaborators, who studies the coexistence of multiple clinical conditions or diseases, such as multimorbidity or comorbidity, which are important topics in clinical medicine and public health [57, 27]. Mechanisms underlying multimorbidity or comorbidity are complex, they may share some direct causation, and associated risk factors like aging and socioeconomic deprivation [44, 60, 54], but the same factors may play different roles in different diseases. The expert hopes to analyze the influencing factors of different diseases and compare those diverse yet correlated diseases or clinical outcomes aiming to gain insights into the development of diseases and improve prognosis.
@@
Following the design study methodology  [52] and the abstraction according to nested model [43], we hosted one in-person meeting in the beginning (July 2022) and several follow-up online discussions (roughly every 2–3 months depending on the progress) with three clinical medicine and public health experts from the university that the first/corresponding author affiliates to (one of them is a co-author who is now at a different university). First, we abstracted the domain problem into two data analysis stages as shown in Fig. 1 that explains our workflow.
@@
    Allow intuitive comparison of causal relationships for different outcomes. Different diseases may share the same influencing factors, and distinguishing the shared and different influencing factors can help medical experts to rationally judge and intervene in specific influencing factors, and improve prognosis.
@@ 3.1 Causal Discovery
Various causal discovery methods are available [62] but none is currently available as an all-around winner. Therefore, we aim to use recent causal discovery methods that inherently handle mixed-type data with categorical and continuous variables and could provide complementing results so that the user can combine them to get a good causal graph. Three causal discovery methods are adapted into our system: the classic Peter-Clark (PC) [55] algorithm, which is known to provide a good causality skeleton, the DAG-GNN method [67] as a representative modern continuous optimization method that can handle mixed-type data, and the hybrid causal learning method (HCM) [40], specifically designed for mixed-type data. More details of the methods are covered in the supplemental material.
@@
The PC [55] algorithm is a well-known and frequently used causal discovery method that serves as the backbone approach. The method starts with a fully connected undirected graph and identifies a skeleton by evaluating the conditional independence of each variable pair. Next, the edges are oriented by identifying colliders/v-structures. The algorithm then orients additional edges in partially directed paths which would otherwise form more v-structures.
@@
The original combinatorial optimization problem of causal discovery can be transformed into a continuous optimization problem [70]. DAG-GNN [67] uses a generative graph neural network based on variational autoencoders with neural network functions and evidence lower bound to resolve the problem. It extends the NO TEARS [70] method for nonlinear structural equation models and supports both discrete and continuous variables.
@@
The HCM is a recent advancement for mixed-type data. It addresses the mixed-type data problem by formulating a mixed structure equation model [40] that treats continuous and categorical variables separately. The model is learned with skeleton learning using the PC-stable algorithm [18] with a new mixed-type conditional independence test; then the method finds causal graph directions via greedy search with a mixed information criterion; and the graph is further pruned using the conditional independence test.
@@ 3.2 Evaluation with Benchmark Datasets
Moreover, with our visualization, patterns can be found that are otherwise difficult with numerical measures only. For example, the coexistence of edges discovered by different methods suggests that the possibility of a true edge is high. In both benchmark data, edges shared by all three methods are all true edges. Therefore, the combination of these algorithms provides users with important additional information for creating a better causal graph than using any of them alone.
@@ 3.3 Comparative Progressive Visualization of Causal Discovery Algorithms
Our visualization method for single-outcome causal graph addresses the following three issues: (1) visually encodes information of a causal graph (R1); (2) supports the visual comparison of results generated by the algorithms in Section 10.1 (R2); (3) visualizes the progress of the algorithms with user-controllable pausing and resuming (R3).
@@ 3.3.1 Visual Encodings of Causal Graphs
An efficient layered method [^3] https://github.com/dagrejs/dagre/wiki that takes advantage of several graph drawing works [21, 35, 8, 11] is used as the layout of the causal graph. In this layout, nodes are arranged top-down with decreasing out-degrees, and the cause-outcome relationship is inherently encoded by the vertical locations of two variables, i.e., the direction of causality naturally has to be from the upper node to the lower node of the two nodes connected by an edge. However, results of causal discovery methods sometimes include reversed causalities with the layered layout, i.e., cause nodes are below effect nodes. Such potentially self-contradictory cases require the attention of users visually. Arrows are commonly used to encode directional properties, so we choose to highlight the reversed causality using an edge with an arrow heading from the cause to the effect and keep other edges with top-down causal directions undirected to reduce clutter.
Color and shape—known as effective visual channels for categorical variables [42]—are used to distinguish outcome and other nodes. Edge thickness, a frequently used channel for weight encoding in a weighted node-link diagram [2], encodes the magnitude of the causal effect (Section 3.4). Note that the causal effect is signed, and we use the edge style to encode the sign: solid lines for positive and dashed lines for negative effect values, respectively. Logarithmic transformation is applied to adapt the large range of the effect value.
Example visualizations are shown in Visual Analysis of Multi-outcome Causal Graphs(1),  Fig. 9(g) and  Fig. 6(4). The node encoding uses jagged circles in orange for the outcome variable, and regular circles in blue for other variables, and the name of each variable is drawn at the center of the circle. The two small widgets (in blue) attached to the top and bottom of a node are used for adding new edges.
@@ 3.3.3 Progressive Visualization
Diagnostic measures of DAG-GNN are visualized as line charts in an epoch-loss view shown at the bottom of Visual Analysis of Multi-outcome Causal Graphs(f), Fig. 9(g), and  Fig. 6(4). Specifically, the measurements — ELBO (Evidence Lower Bound) loss (blue), NLL (Negative Log-Likelihood) loss (green), and MSE (Mean Squared Error) loss (orange) —provide interpretability of the method and evidence for users to determine when to pause or stop the algorithm. Users can pause the process at any time to examine the current result and stop or restart the process depending on whether the result is satisfactory or not. Typically, the losses drop quickly after dozens of epochs and become stable and a similar happens to the associated causal graph result. To this end, our visualization provides more insights into the DAG-GNN method that was previously impossible.

### v2

@@ Abstract
The overall process of the visual analysis of multi-outcome causal graphs of a health research dataset [15]. Analysts use (1) the Single Graph View to analyze and edit the causal graph (a–c) of a single outcome aided by a set of new visualization techniques (d–f) to leverage the synergy of various causal discovery methods. The fine-tuned causal graphs of single outcomes are stored in (2) the History View (g). Next, causal graphs of interest are selected for visual comparison in (3) the Multi-outcome Graphs Comparison View (h, i), supported by our new graph layout and visual mappings (i).
@@ Introduction
In medicine, experts also seek to discover causalities between potential risk factors and a specific disease, aiming to explore factors influencing the development of a particular disease and eventually intervene in the prognosis. Except for the typical case of studying the causality of a single disease (a single outcome variable), understanding the similarities and differences between different yet related causal relationships among different diseases, i.e., multiple outcome variables, is also important. For example, a study into the co-occurrence of more than one chronic condition in one person (multimorbidity [28]) and the combined effect of multiple conditions in one person (comorbidity [60]) is beneficial for proper interventions and treatments. Therefore, it is necessary to perform analysis and comparisons for causality of multiple outcomes. Unfortunately, most previous visual analysis works intend to address problems with only one particular outcome.
In this paper, we study the causalities of multiple outcome variables that have not been considered in visual analytics. Based on a requirement analysis in close collaboration with medical experts, we propose a visual analysis framework as shown in Visual Analysis of Multi-outcome Causal Graphs. The framework consists of two stages: the discovery and interactive fine-tuning of a causal graph of a single outcome (Visual Analysis of Multi-outcome Causal Graphs(1)), and the comparative analysis of causal graphs of multiple outcomes (Visual Analysis of Multi-outcome Causal Graphs(3)).
Causal graphs in the first stage are calculated using three causal discovery algorithms: a classic causal discovery method that delivers a basic graph, and two state-of-the-art methods that provide analysts with more suggestions. The resulting graphs are visualized progressively as node-link diagrams with a superposed comparative visualization for easy detection of differences. The underlying machine learning methods can be interactively controlled, e.g., by pausing and resuming, to reduce computation time. Users then interactively fine-tune the single outcome causal graph with reference to multiple views of multivariate data visualizations and with their domain knowledge.
@@
We evaluate our method with numerical measurements, a case study, and an expert user study with medical experts. The effectiveness of the analysis and visualization of single outcome graphs is demonstrated using benchmark datasets; the multi-outcome comparison design is evaluated with stress measurements. The overall visual analysis method is evaluated through a case study of the analysis of diseases that are usually seen in multimorbidity/comorbidity using two widely accepted health research datasets. Feedback from expert users indicates that our method is useful for supporting causal reasoning of single and multiple outcomes and provides insights with the comparison of multi-outcome causal graphs that are impossible with existing methods.
@@
Please note that our method aims to address a domain problem in healthcare but is also general enough to aid causal analysis of single and multiple outcomes. The source code of our visual analysis system is publicly available at: https://github.com/mengjiefan/multi_outcome.
@@ 1 Related Work
Our work is related to three areas of research: causal discovery, visualization and visual analysis of causality, and comparative visualization.
@@ 1.1 Causal Discovery
Combinatoric methods search for DAGs satisfying criteria within the aforementioned four categories. Popular combinatoric methods include the Peter-Clark (PC) algorithm [55], Greedy Equivalence Search (GES) [16], satisfiability (SAT) solver [32], and LinGAM [53].
In contrast, continuous methods rely on continuous optimizations that are typically used in neural networks. NO TEARS (Non-combinatoric Optimization via Trace Exponential Augmented Lagrangian Structure learning) is considered the first to formulate the combinatorial graph search method as a continuous optimization problem [70]. A further improvement of the method that is faster and more accurate than NO TEARS is available [9]. DAG-GNN [67] is a graph neural network approach to solve the continuous optimization problem. Methods focused on categorical data are also available with continuous optimizations using interventional or reinforcement learning [37, 12].
One issue with causal discovery methods is that they typically assume that all variables have the same data type—either categorical or continuous. In real-world datasets, however, data types are mixed, for example, typical survey data in medicine includes continuous biological measures like body height, weight, and blood pressure, and binary/categorical attributes like gender and stages of diseases. Few, dedicated methods are available for mixed-type datasets [3, 4, 40, 14], while some others, for example, DAG-GNN [67], can handle such data inherently. A latent-PC algorithm is available for multi-dimensional mixed-type data that is typical in medical research [14].
@@ 1.2 Visualization of Causal Graphs and Visual Analysis of Causal Reasoning
The visualization of causal graphs is studied in several works [6, 66, 61, 19]. Different visual mappings of the graph, e.g., adjacent matrix and node-link diagram are compared in a user study [61]. For node-link diagrams, different graph layouts are also compared with user studies [6, 61]. Layered graph layouts are used for showing large causal graphs [66] or comparing different causal discovery results [19].
Visual analysis methods for causal reasoning are available to aid the understanding of causalities and support interactive decision-making. An interactive causal reasoning interface allows visually editing and verifying causal relationships with multiple linked views of the causal graph, and statistical parameters [63]. A more comprehensive extension of the method is available with the support of multidimensional data visualizations, a visual design of the causal graph with layered graph layout, and the capability of handling mixed-type data [64]. An exploratory causal analysis approach introduces the uncertainty-aware visualization of causal graphs and provides interactive means for exploring, validating, and applying causal relations for decision-making [66].
Some works focus on the understanding of causal inferences and counterfactuals [38]. A multi-view visualization tool is available to support iterative causal inference [26]. Counterfactuals are used to improve the visualization of relationships between variables [36] and causal inferences [10]. Visual analytics of time-dependent causality can be applied to event sequences with a user-modifiable Granger causality model [34]. Causal hypotheses of time-dependent events can be formulated and tested with time delay in a visual analysis method [65].
@@ 1.3 Comparative Visualization
It is a common yet challenging task to design methods supporting comparison in visualization [24]. Four considerations are presented for comparative visualization designing and evaluation [24]. A general taxonomy of visual comparisons classifies visual designs into three basic categories—juxtaposition, superposition, and explicit encodings [25]. These categories should be combined for effective comparative visualization tools and tasks. A qualitative user study shows that interactive techniques are important to visual comparison tasks [58]. A combined encoding of colors and positions enables analysts to easily view differences and changes in the comparison of task-driven topic models [1].
Research efforts have been made for the visual comparison of graphs. Visual comparison of two graphs is achieved by visualizing their similarities and differences in a merged graph [5]. Interactive graph matching [29] addresses the visual comparison of graphs and clustered graphs. Weighted graph comparison is explored in brain connectivity analysis [2], where the effectiveness of node-link diagrams and adjacency matrices is compared. A map-based abstraction of node-aligned graphs on a triangle mesh facilitates a clutter-reduced comparison of undirected graphs [33]. Layout optimization, graph embedding, and clustering and grouping algorithms are used for the comparison of two temporal graph datasets [69].
@@ 2 Requirement Analysis and Method Overview
Our work is a long-term collaboration with medical experts, which started in 2022. The work was initially motivated by the needs of one of our collaborators who studies the coexistence of multiple clinical conditions or diseases, such as multimorbidity or comorbidity, which are important topics in clinical medicine and public health [57, 27]. Mechanisms underlying multimorbidity or comorbidity are complex, they may share some direct causation, and associated risk factors like aging and socioeconomic deprivation [44, 60, 54], but the same factors may play different roles in different diseases. The expert hopes to analyze the influencing factors of different diseases and compare those diverse yet correlated diseases or clinical outcomes aiming to gain insights into the development of diseases and improve prognosis.
@@
Following the design study methodology [52] and the abstraction according to a nested model for visualization design [43], we hosted one in-person meeting in the beginning (July 2022) and several follow-up online discussions (roughly every 2–3 months depending on the progress) with three clinical medicine and public health experts from the university that the first/corresponding author affiliates to (one of them is a co-author who is now at a different university). First, we abstracted the domain problem into two data analysis stages as shown in Fig. 1 that explains our workflow.
@@
    Allow intuitive comparison of causal relationships for different outcomes. Different diseases may share the same influencing factors, and distinguishing the shared and different influencing factors can help medical experts rationally judge and intervene in specific influencing factors, and improve prognosis.
@@ 3.1 Causal Discovery
There are various causal discovery methods [62], but none is currently available as an all-around winner. Therefore, we aim to use recent causal discovery methods that inherently handle mixed-type data with categorical and continuous variables and could provide complementing results so that the user can combine them to get a good causal graph. Three causal discovery methods are adapted into our system: the classic Peter-Clark (PC) [55] algorithm, which is known to provide a good causality skeleton, the DAG-GNN method [67] as a representative modern continuous optimization method that can handle mixed-type data, and the hybrid causal learning method (HCM) [40], specifically designed for mixed-type data. More details of the methods are covered in the supplemental material.
@@
The PC [55] algorithm is a well-known and frequently used causal discovery method that serves as the backbone approach. The method starts with a fully connected undirected graph and identifies a skeleton by evaluating the conditional independence of each variable pair. Next, the edges are oriented by identifying colliders/v-structures. The algorithm then orients additional edges in partially directed paths that would otherwise form more v-structures.
@@
The original combinatorial optimization problem of causal discovery can be transformed into a continuous optimization problem [70]. DAG-GNN [67] uses a generative graph neural network based on variational autoencoders with neural network functions and evidence lower bound to solve the problem. It extends the NO TEARS [70] method for nonlinear structural equation models and supports both discrete and continuous variables.
@@
HCM is a recent advancement for mixed-type data. It addresses the mixed-type data problem by formulating a mixed structure equation model [40] that treats continuous and categorical variables separately. The model is learned with skeleton learning using the PC-stable algorithm [18] with a new mixed-type conditional independence test; then the method finds causal graph directions via greedy search with a mixed information criterion; and the graph is further pruned using the conditional independence test.
@@ 3.2 Evaluation with Benchmark Datasets
Moreover, with our visualization, patterns can be found that are otherwise difficult with numerical measures only. For example, the coexistence of edges discovered by different methods suggests that the possibility of a true edge is high. In both benchmark datasets, edges shared by all three methods are all true edges. Therefore, the combination of these algorithms provides users with important additional information for creating a better causal graph than using any of them alone.
@@ 3.3 Comparative Progressive Visualization of Causal Discovery Algorithms
Our visualization method for single-outcome causal graph addresses the following three issues: (1) it visually encodes information of a causal graph (R1), (2) supports the visual comparison of results generated by the algorithms in Section 10.1 (R2), and (3) visualizes the progress of the algorithms with user-controllable pausing and resuming (R3).
@@ 3.3.1 Visual Encodings of Causal Graphs
An efficient layered method [^3] https://github.com/dagrejs/dagre/wiki that takes advantage of several graph drawing works [21, 35, 8, 11] is used for the layout of the causal graph. In this layout, nodes are arranged top-down with decreasing out-degrees, and the cause-outcome relationship is inherently encoded by the vertical locations of two variables, i.e., the direction of causality naturally is from the upper node to the lower node of the two nodes connected by an edge. However, results of causal discovery methods sometimes include reversed causalities with the layered layout, i.e., cause nodes are below effect nodes. Such potentially self-contradictory cases require user attention. Arrows are commonly used to encode directional properties, so we choose to highlight the reversed causality using an edge with an arrow heading from the cause to the effect and keep other edges with top-down causal directions undirected to reduce clutter.
Color and shape—known as effective visual channels for categorical variables [42]—are used to distinguish outcome and other nodes. Edge thickness, a frequently used channel for weight encoding in a weighted node-link diagram [2], shows the magnitude of the causal effect (Section 3.4). Note that the causal effect is signed, and we use the edge style to encode the sign: solid lines for positive and dashed lines for negative effect values, respectively. Logarithmic transformation is applied to adapt the large range of the effect value.
Example visualizations are shown in Visual Analysis of Multi-outcome Causal Graphs(1), Fig. 9(g), and Fig. 6(4). The node encoding uses jagged circles in orange for the outcome variable, and regular circles in blue for other variables, and the name of each variable is drawn at the center of the circle. The two small widgets (in blue) attached to the top and bottom of a node are used for adding new edges.
@@ 3.3.3 Progressive Visualization
Diagnostic measures of DAG-GNN are visualized as line charts in an epoch-loss view shown at the bottom of Visual Analysis of Multi-outcome Causal Graphs(f), Fig. 9(g), and Fig. 6(4). Specifically, the measurements—ELBO (Evidence Lower Bound) loss (blue), NLL (Negative Log-Likelihood) loss (green), and MSE (Mean Squared Error) loss (orange)—provide interpretability of the method and evidence for users to determine when to pause or stop the algorithm. Users can pause the process at any time to examine the current result and stop or restart the process depending on whether the result is satisfactory or not. Typically, the losses drop quickly after a few dozen epochs and become stable, and similar is true for the associated causal graph result. Therefore, our visualization provides more insights into the DAG-GNN method than what was previously possible.